OpenStudy (anonymous):

please help me solve 6z / 5x = 30 10x+20y=40 4x+7y=28

OpenStudy (anonymous):

For 10x+20y=40 10x+0=40 10x=40 x=4 0 + 20y = 40 20y = 40 y=2

OpenStudy (anonymous):

For 4x + 7y =28 4x + 0 = 28 4x=28 x=7 0 + 7y = 28 7y= 28 y=4

OpenStudy (anonymous):

do you know how to do the first one?

OpenStudy (anonymous):

not really I'll see what I can do

OpenStudy (anonymous):

okay thank you